OpenStudy (anonymous):

Let u = <-6, 3>, v = <-1, -6>. Find 4u + 2v.

OpenStudy (anonymous):

ok so you are going to start by plugging in -6 for u and -1 for v which should get you 4(-6)+2(-1)=-26. then plug in 3 and -6 to get 4(3)+2(-6)=0 then plug in -6 and -6 which gets 4(-6)+2(-6)=-36. finally plug in 3 and -1 to get 4(3)+2(-1)=10

OpenStudy (anonymous):

thank you so much! could you help me with a few more?

OpenStudy (anonymous):

of course

OpenStudy (anonymous):

Determine whether the vectors u and v are parallel, orthogonal, or neither. u = <8, 4>, v = <10, 7>

OpenStudy (anonymous):

they are orthogonal!

OpenStudy (anonymous):

Given that P = (7, 16) and Q = (-1, 7), find the component form and magnitude of qp